# Equality Saturation for Optimizing High-Level Julia IR

### v1

@@ Abstract.
Compilers are indispensable for transforming code written in high-level languages into performant machine code, but their general-purpose optimizations sometimes fall short. Domain experts might be aware of certain optimizations that the compiler is unable to apply or that are only valid in a particular domain. We have developed a system that allows domain experts to express rewrite rules to optimize code in the Julia programming language. Our system builds on e-graphs and equality saturation. It can apply optimizations in the presence of control flow and side effects. As Julia uses multiple dispatch, we allow users to constrain rewrite rules by argument types, and propagate type information through the e-graph representation. We propose an ILP formulation for optimal e-graph extraction taking into account dominance properties for code reuse and introduce CFG skeleton relaxation to rewrite calls to pure functions as well as those with side effects. Use cases demonstrate that our system can perform rewrites on high-level, domain-specific code, as well as on lower-level code such as Julia’s broadcasting mechanism. Finally, we analyze the required compilation time.
@@ 1. Introduction
Optimizing compilers have become indispensable as programming languages rely on them to unravel high-level abstractions into lean and performant code (Lattner et al., 2020; Hoste and Eeckhout, 2008; Lattner and Adve, 2004).
@@
General-purpose compiler middle-ends typically perform general-purpose optimizations such as common subexpression elimination (Muchnick, 1998). Compilers for domain-specific languages perform domain-specific optimizations such as algebraic optimizations (Lattner et al., 2021; Muchnick, 1998). The front-ends of higher-level language compilers can be customized to generate IR that is already optimized for specific targets, such as GPUs, and then fed to the back-ends (Besard et al., 2019). Many compilers can also be configured to enable context-specific optimizations, e.g., to sacrifice precision for performance.
However, none of those existing optimization pipelines is well suited to support domain-specific and context-specific optimization of code written in a high-level scientific programming language that is used across many domains, such as Julia (Bezanson et al., 2012). In such a language, software libraries build on the generic language infrastructure to provide high-level domain-specific APIs to application developers.
@@
Ideally, those application developers should be able to enjoy domain-specific optimizations for free, i.e., without having to manually tune or rewrite their code, and without facing restrictions on their freedom to exploit the language’s rapid prototyping features. Furthermore, it should be easy for application developers to specify or select which context-specific optimization they want to see enabled. In addition, when multiple libraries are being reused and composed in an application, be it top-level libraries that provide orthogonal functionality or higher-abstraction-level libraries that provide more abstract APIs on top of lower-level code, the optimization opportunities coming with those libraries should be composable. They should be composable with each other and with the general-purpose and target-specific optimizations that are provided in the main compiler flow and in the target-specific extensions thereof.
This inevitably involves solving a phase-ordering problem because some optimizations create opportunities for further optimizations, while others are incompatible and rule each other out. Finally, to obtain a thriving ecosystem of domain libraries, the development of such libraries and the corresponding support for domain-specific optimization should not require the involvement of compiler experts.
It then follows that a generic compiler infrastructure is needed that enables developers of domain-specific libraries to specify domain-specific and context-specific optimization opportunities at an abstraction level similar to that of their libraries’ APIs, and that ensures sufficient composability.
@@
With our research, we aim to provide that generic compiler infrastructure. In this paper, we present a system that allows programmers, not necessarily compiler developers, to express their domain-specific knowledge as rewrite rules in Julia. Our work is based in part on ideas developed for the Cranelift compiler (Fallin, 2023), which in turn uses e-graphs and equality saturation, first introduced by Nelson (Nelson, 1980) and Tate et al. (Tate et al., 2009), respectively. Equality saturation is a rewrite technique that elegantly deals with the phase ordering problem by representing the potential results of rewrites while still representing the original program as well.
@@
When writing and executing code, users should not take these rewrite rules into account anymore, but instead can focus on writing simple and readable code that maps well on their mental model of the computations. The compiler then automatically evaluates all relevant rewrites and picks the optimal program. By building on e-graphs and equality saturation, the phase ordering problem is no longer an issue. Rewrite rules are applied without removing information from the e-graph, which means that application of a rewrite cannot prevent other rules from firing.
This paper’s contributions are as follows:
the first use of equality saturation in a high-level, dynamically typed programming language optimizer, including for code containing control flow;
CFG skeleton relaxation to enable rewriting on functions with side-effects;
a novel ILP formulation to enable value reuse in (rewritten) expressions;
a novel e-class analysis to track type information for type-constraint rewrite rules;
a demonstration on a number of examples;
an analysis of the required compilation time.
@@ 2. Background
The e-graph data structure compactly represents a congruence relation on different expression trees (Nelson, 1980; Willsey et al., 2021). It consists of a collection of e-classes that can each contain multiple e-nodes representing expression trees that are equivalent to each other according to some user-defined equivalence relation. Each e-node can have multiple children, represented by e-classes. As an example, the left side of Figure 1 shows the e-graph representation for the expression . In case represents an integer, a potential equivalent expression for is . We can encode this equivalence by adding a new e-node to the appropriate e-class, as shown on the right side of Figure 1. Because the constant expression 1 is not equivalent to any of the e-classes already present, it is added to a newly created e-class. The strength of e-graphs lies in the fact that both equivalent representations are encoded in the graph at the same time.
@@
Besides in theorem solvers (de Moura and Bjørner, 2007; Nelson, 1980; De Moura and Bjørner, 2008), e-graphs are often used for equality saturation (Willsey et al., 2021; Cheli, 2021b; Tate et al., 2009). Equality saturation is a term rewriting technique that applies rewrites not by overwriting the original term but by adding the rewritten term to an e-class that represents the original term. By building on e-graphs for term rewriting, it is possible to represent a much larger amount of rewritten, equivalent expressions without the exponential explosion that can occur with naive approaches. The advantage of not discarding expressions when they get rewritten is that the ordering in which rewrites are applied cannot prevent certain rewrites from firing.
@@
Once no more new rules can be applied or some timeout is reached during equality saturation, one typically wants to determine the optimal expression contained in the e-graph (Willsey et al., 2021; Goharshady et al., 2024). In its simplest form, extraction starts from a root e-class where one e-node needs to be picked to be extracted. Recursively, all the child e-classes of the picked e-node need to be extracted as well. The end result is a directed, connected graph of e-nodes.
In many use cases, especially in program optimization, the extracted graph cannot contain cycles. In such cases, more refined extraction techniques are needed. One such technique, which considers additional constraints from our problem domain, is discussed in Section 3.6.
@@ 3. A Rewrite-Based Compiler Middle-End
Our system allows programmers to specify custom rewrite rules. These rules are then applied on an e-graph representing the code of a function. Once the rules are applied and the e-graph is saturated, the compiler will try to extract the optimal code embedded within. Some example rules are
@@
Rules have a left-hand and right-hand side consisting of a Julia code expression, where the ~ prefix is used to indicate unbound variables. The :: notation is used to specify the types of variables for which the rule can be applied. sin, cos, and translate are the names of called functions. A rewrite rule can be applied whenever the pattern of function calls, literals, and variables on its left-hand side matches a code fragment.
@@
As shown, the two sides of a rewrite rule are specified as regular Julia code expressions, with the addition of unbound, optionally typed, variables. This syntax is able to cover all rewrites where a single expression tree is replaced with a new expression tree. This syntax does not allow rules matching multi-patterns (Jia et al., 2019; Yang et al., 2021). These patterns can match multiple expression trees and rewrite them each, potentially reusing variables in multiple output expressions.
3.1. Incorporating Rewrites in the Compiler
On top, the Julia compiler pipeline is shown: Julia code is parsed into an AST, which is lowered to Julia IR, for which Julia Codegen produces LLVM IR, which is further compiled by LLVM Codegen. On the bottom, the figure zooms in on additional stages that are executed on the Julia IR. These are macro expansion and control flow destructuring to yield untyped IR, type inference and constant propagation that produce typed IR, inlining to obtain inlined IR and finally SROA and simplification to yield optimized IR.
@@ 3.1. Incorporating Rewrites in the Compiler
As in most compiler middle-ends, this optimization will operate on an intermediate representation (IR) that supports optimizations while still preserving high-level language-specific information (Lattner et al., 2020; Li et al., 2021; swiftlang, 2012). In our case, this is the Julia compiler’s SSA-based IR (Bezanson et al., 2012).
Figure 2 shows the flow of the compiler. When a function needs to be executed, the just-ahead-of-time flow is invoked on it. We will call this the current function. The compiler first lowers the current function’s AST into an SSA-based IR. On this IR, it applies type inference, constant propagation, function inlining, and a small number of other basic optimizations. From this Julia IR, the compiler then generates LLVM IR, to which many more optimizations are applied by the LLVM compiler before it generates assembly code for the target architecture.
There are two places where it makes sense to apply our rewrite optimizations: right before inlining, and after inlining and optimizations. Before inlining, the current function’s IR still contains function calls as they were written by the programmer. At this stage in the compilation, high-level rewrite rules can typically be matched and applied, i.e., rules involving functions of high-level abstract APIs.
In some cases, it might be useful to also apply rewrites to code in which function calls have already been inlined. After inlining, it is possible that code from different function bodies down the call chain of the current function ends up in its IR. This can uncover additional rewrite opportunities.
@@
Our current prototype implementation performs the rewriting optimization both before and after inlining, but we note that the most interesting rewrites fire before inlining, when high-level API usage has not yet vanished from the code. For this reason, the example snippets of IR in this paper contain code that has not been inlined yet.
@@ 3.2. From IR to E-Graph and Back Again
Top left: a simple Julia function that computes x to the power n and that prints a message to illustrate the impact of side effects. Top right: the corresponding Julia IR that is generated if the power function is generated for two integer arguments. It consists of four basic blocks, one of which includes two phi functions. Bottom left: the e-graph for the expressions in the function consisting of 11 e-classes, including classes for the phi functions and for the side effect. Bottom right: the CFG skeleton of the function, which looks like the CFG itself, and in which each node refers to certain e-classes in the e-graph where the control flow depends on expressions modeled by those e-classes.
To apply rewrites on the Julia IR of a function, we first convert it to an e-graph. For this, we employ the methods introduced in the Cranelift compiler (Fallin, 2023). Each function call is added to an e-graph, while control flow statements (goto, goto if not, phi, and return) are added to an external data structure called the CFG skeleton. Function calls that have side effects are also added to the CFG skeleton. Each statement in the CFG skeleton refers to the e-nodes on which it depends. Figure 3 shows an example Julia function pow together with its corresponding IR, e-graph, and CFG skeleton. In addition to the control flow statements, the call to println has been stored in the CFG skeleton because calling that function produces side effects. In this representation, the phi nodes are leaf nodes. Their dependencies are not included in the e-graph. This implies (i) that the e-graph does not contain sufficient information to reconstruct all control flow of the original code, and (ii) that rewrites including phi-node dependencies are not supported. It is the CFG skeleton that provides the extra information needed to reconstruct valid IR.
@@
It is quite common that part of a function depends on a kwarg, e.g., to perform extra checks or to modify the behavior. See, e.g., the use of the flipkernel kwarg in NNLib.jl.111https://github.com/FluxML/NNlib.jl/blob/81e6cd1843dbd2baf2b7b2efd00e942f4f040a2d/src/fold.jl#L186-L191.
@@
BLAS/LAPACK-like libraries often use characters to indicate the format of an array (transposed or not, upper/lower triangular, etc.). Such checks are then independent of the values calculated beforehand. For example, in the LinearAlgebra.jl222https://github.com/JuliaLang/LinearAlgebra.jl/blob/master/src/bidiag.jl#L1414-L1448. standard library, the costly operation blks = findall(...) is first computed, after which there is a branch based on uplo, followed by a reuse of blks. Another example is CUDA.jl’s CUSPARSE wrappers: Plenty of API calls (which are a prime use case for e-graph based rewriting) occur outside of fmt checks, others occur inside such checks.333https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusparse/generic.jl#L48-L70.
@@
Continuing on the subject of API calls, it is common that certain API uses are conditional on the version of the library. For example, CUDA.jl includes checks such as CUSPARSE.version() v"11.7.2".444https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusparse/generic.jl#L304-L335.
@@
In scalar ‘mathy’ code, this pattern is also common. The absolute cost of operations outside the if is then often not that great, but when it comes to optimizing very high-performance code, it can make a difference. Some examples can be found in the standard float.jl and math.jl libraries.555https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/float.jl#L320-L328666https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/math.jl#L756-L766.
@@
For readability, or to avoid too many methods, imperative type checks are also sometimes used in Julia instead of multiple dispatch. Several examples can be found in the standard libraries.777https://github.com/JuliaGPU/CUDA.jl/blob/860eb88e40053b2709ef949f2eaf593c59bcecf1/lib/cusolver/linalg.jl#L35-L48, https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/abstractarray.jl#L2485-L2497, https://github.com/JuliaLang/julia/blob/9118ea7565feae13d5b47654a3c245c0df36e753/base/intfuncs.jl#L592-L606.
@@
Like MLIR and the Swift Intermediate Language (SIL) (swiftlang, 2012; MLIR, 2019), Cranelift’s SSA IR format uses basic block arguments, which are bound by the branch operations in a block’s predecessors, to represent control-dependent data flow. Julia’s SSA IR, by contrast, uses phi-nodes. Basic block arguments are more convenient for materializing statements in the e-graph because all information regarding out-of-block uses of expressions that need to be materialized within a block are then explicitly represented within that block itself.
For this reason, our code representation differs from existing uses of e-graphs and CFG skeletons in that it contains both phi-nodes and basic block arguments. Take, for example, the two goto #2 statements in the Julia IR in Figure 3. Since the destination block (#2) contains two phi nodes, the goto statements in the CFG skeleton pass the respective values as arguments. The e-graph itself still stores phi node values, but these are now handled as basic block arguments. During conversion from IR to e-graph and back from e-graph to IR, a conversion between phi nodes and basic block arguments is therefore performed.

### v2

@@ Abstract.
Compilers are indispensable for transforming code written in high-level languages into performant machine code, but their general-purpose optimizations sometimes fall short. Domain experts might be aware of optimizations that the compiler is unable to apply or that are only valid in a particular domain. We have developed a system that allows domain experts to express rewrite rules to optimize code in the Julia programming language. Our system builds on e-graphs and equality saturation. It can apply optimizations in the presence of control flow and side effects. As Julia uses multiple dispatch, we allow users to constrain rewrite rules by argument types, and propagate type information through the e-graph representation. We propose an ILP formulation for optimal e-graph extraction that exploits opportunities for code reuse and introduce CFG skeleton relaxation to rewrite calls to pure functions as well as those with side effects. Use cases demonstrate that our system can perform rewrites on high-level, domain-specific code, as well as on lower-level code such as Julia’s broadcasting mechanism. We analyze the required compilation time and the performance impact of these rewrites.
@@ 1. Introduction
Optimizing compilers have become indispensable as programming languages rely on them to unravel high-level abstractions into lean and performant code (Lattner et al., 2021; Hoste and Eeckhout, 2008; Lattner and Adve, 2004). General-purpose compiler middle-ends typically perform general-purpose optimizations such as common subexpression elimination (Muchnick, 1998). Compilers for domain-specific languages perform domain-specific optimizations such as algebraic optimizations (Lattner et al., 2021; Muchnick, 1998). The front-ends of higher-level language compilers can be customized to generate IR that is already optimized for specific targets, such as GPUs, and then fed to the back-ends (Besard et al., 2019). Many compilers can also be configured to enable context-specific optimizations, e.g., to sacrifice precision for performance. However, none of those existing optimization pipelines are well suited to support domain-specific and context-specific optimization of code written in a high-level scientific programming language that is used across many domains, such as Julia (Bezanson et al., 2012). In such a language, software libraries build on the generic language infrastructure to provide high-level domain-specific APIs to application developers.
@@
Ideally, those application developers should be able to enjoy domain-specific optimizations for free, i.e., without having to manually tune or rewrite their code, and without facing restrictions on their freedom to exploit the language’s rapid prototyping features. Furthermore, it should be easy for application developers to specify or select the context-specific optimization they want enabled. In addition, when multiple libraries are being reused and composed in an application, be it top-level libraries that provide orthogonal functionality or higher-abstraction-level libraries that provide more abstract APIs on top of lower-level code, the optimization opportunities coming with those libraries should be composable. They should be composable with each other and with the general-purpose and target-specific optimizations that are provided in the main compiler flow and in the target-specific extensions thereof. This inevitably involves solving a phase-ordering problem because some optimizations create opportunities for further optimizations, while others are incompatible and rule each other out. Finally, to obtain a thriving ecosystem of domain libraries, the development of such libraries and the corresponding support for domain-specific optimization should not require the involvement of compiler experts. It then follows that a generic compiler infrastructure is needed that enables developers of domain-specific libraries to specify domain-specific and context-specific optimization opportunities at an abstraction level similar to that of their libraries’ APIs, and that ensures sufficient composability.
@@
With our research, we aim to provide that generic compiler infrastructure. In this paper, we present a system that allows programmers, not necessarily compiler developers, to express their domain-specific knowledge as rewrite rules in Julia. Our work is based in part on ideas developed for the Cranelift compiler (Fallin, 2023), which in turn uses e-graphs and equality saturation, first introduced by Nelson (Nelson, 1980) and Tate et al. (Tate et al., 2009), respectively. Equality saturation, or EqSat, is a rewrite technique that elegantly deals with the phase-ordering problem by representing the potential results of rewrites while still representing the original program as well.
@@
When writing and executing code, users should not take these rewrite rules into account anymore, but instead can focus on writing simple and readable code that maps well onto their mental model of the computations. The compiler then automatically evaluates all relevant rewrites and picks the optimal program. By building on e-graphs and EqSat, the phase-ordering problem is no longer an issue. Rewrite rules are applied without removing information from the e-graph, which means that application of a rewrite cannot prevent other rules from firing.
The contributions in this paper are as follows: (i) the first use of EqSat in a high-level, dynamically typed programming language optimizer, including for code containing control flow; (ii) CFG skeleton relaxation to enable rewriting on functions with side effects; (iii) a novel ILP formulation to enable value reuse in (rewritten) expressions; (iv) a novel e-class analysis to track type information for type-constrained rewrite rules; (v) a demonstration on a number of examples; (iv) an analysis of the required compilation time; (vi) an analysis of the potential performance gains achievable.
This paper is structured as follows. Section 2 provides background. Section 3 presents our rewriting system. Section 4 presents its capabilities on use cases. Section 5 analyzes its compilation time, and Section 6 demonstrates obtainable performance optimizations. Sections 7 and 8 discuss limitations and related work. Finally, Section 9 draws conclusions.
@@ 2. Background
E-Graphs. An e-graph compactly represents a congruence relation on different expression trees (Nelson, 1980; Willsey et al., 2021). It consists of a collection of e-classes that can each contain multiple e-nodes representing expression trees that are equivalent to each other according to a user-defined equivalence relation. Each e-node can have multiple children, represented by e-classes. As an example, the left side of Fig. 1 shows the e-graph for the expression . In case represents an integer, a potential equivalent expression for is . We can encode this equivalence by adding a new e-node to the appropriate e-class, as shown on the right side of Fig. 1. Because the constant expression 1 is not equivalent to any of the e-classes already present, it is added to a newly created e-class. The strength of e-graphs lies in the fact that both equivalent representations are encoded in the graph at the same time.
@@
Besides in theorem solvers (de Moura and Bjørner, 2007; Nelson, 1980; De Moura and Bjørner, 2008), e-graphs are often used for EqSat (Willsey et al., 2021; Cheli, 2021b; Tate et al., 2009). EqSat is a term rewriting technique that applies rewrites not by overwriting the original term but by adding the rewritten term to an e-class that represents the original term. By building on e-graphs for term rewriting, it is possible to represent a much larger number of rewritten, equivalent expressions without the exponential explosion that can occur with naive approaches. The advantage of not discarding expressions when they get rewritten is that the ordering in which rewrites are applied cannot prevent certain rewrites from firing. Central to EqSat is the rebuilding procedure, which is responsible for maintaining the e-graph invariants. When rewriting uncovers two different expressions to be equivalent, all expressions containing these equivalent expressions as subexpressions need to be verified, as these potentially also need to become equivalent now.
@@
Once no more new rules can be applied or some timeout is reached during EqSat, one typically wants to determine the optimal expression contained in the e-graph (Willsey et al., 2021; Goharshady et al., 2024). In its simplest form, extraction starts from a root e-class where one e-node needs to be picked to be extracted. Recursively, all the child e-classes of the picked e-node need to be extracted as well. The end result is a directed, connected graph of e-nodes.
In many use cases, especially in program optimization, the extracted graph cannot contain cycles. In such cases, more refined extraction techniques are needed. Ours will be discussed in Section 3.5.
Julia
Julia is a general-purpose programming language that is primarily used for scientific computing. Being a dynamically typed language, functions can be defined for unconstrained argument types. A function can also be redefined for specific argument types (and counts) to express that it should behave differently for that combination of types. The different behaviors, i.e., the definitions of a function for different types and counts of arguments, are called methods in Julia. In essence, a function is an identifier, such as f or print, that it shares with all its methods. When a function call is to be executed, the Julia run-time system will look up and dispatch the function’s method that best matches the types and count of all the call’s run-time arguments. This is called multiple dispatch.
In Julia, operators are syntactic sugar for functions. Even basic functionality like array indexing can be customized with methods for different types of arrays and indices. A method can, e.g., redefine how elements in a Symmetric matrix are accessed, avoiding the need to store symmetric elements twice. Methods that implement more abstract computations on matrices by accessing those matrices’ elements through index operations will automatically benefit from that custom access method when they are invoked on a Symmetric matrix. Because all operators are functions, Julia expressions correspond by and large to trees of function calls.
On top, the Julia compiler pipeline is shown: Julia code is parsed into an AST, which is lowered to Julia IR, for which Julia Codegen produces LLVM IR, which is further compiled by LLVM Codegen. On the bottom, the figure zooms in on additional stages that are executed on the Julia IR. These are macro expansion and control flow destructuring to yield untyped IR, type inference and constant propagation that produce typed IR, inlining to obtain inlined IR and finally scalar replacement of aggregates (SROA) and simplification to yield optimized IR.
Fig. 2 shows Julia’s just-ahead-of-time compiler flow. At run time, when a method needs to be dispatched and executed, this flow is invoked on it. We will call this the current method. The compiler first lowers the method’s AST into an SSA-based IR. Because all operators are functions, that IR code consists mainly of control flow, phi-nodes, literals, assignments to SSA values, and function calls. It hence directly reflects all expressions from the source code, including all called functions and operators, the main difference being that each sub-expression result is assigned to its own SSA value in the IR. On this IR, the compiler applies type inference based on the run-time argument types, constant propagation, method inlining, and some other basic optimizations such as scalar replacement of aggregates (SROA). From this Julia IR, the compiler then generates LLVM IR, to which many more optimizations are applied by the LLVM compiler before it generates assembly code for the target architecture.
Cranelift and Acyclic E-Graphs
Cranelift (Fallin, 2023), a compiler back-end mostly used to generate code from Rust and WebAssembly, is the first production compiler to use e-graphs for code optimization. As Cranelift aims at JIT compilation, where compilation time is critical, they have introduced acyclic e-graphs (ægraphs). Like regular e-graphs, these represent multiple equivalent expressions in e-classes. But, while regular EqSat fully propagates the effects of introduced equalities, this is not the case with ægraphs. Equalities are greedily applied, and e-graph invariants are not fully restored. This means that some rewrites are potentially not discovered, but the compilation time is shorter and more predictable.
More relevant to our work is the introduction of the CFG skeleton, a data structure kept alongside the e-graph to keep track of control flow instructions and instructions with side effects, two constructs that do not fit the EqSat paradigm for optimizing expressions. Its workings and how we leverage this technique in our work will be explained in Section 3.1. Because e-graphs in Cranelift are used in the compiler back-end, they model expressions extracted from the IR after inlining, which can be relatively large, but mainly contain simple, low-level instructions, many of which are pure. Cranelifts’ e-graphs and its use of the CFG skeleton are hence designed to conserve all control flow and side effects, instead focusing on the rewriting of pure expressions in between control flow and operations with side effects.
@@ 3. A Rewrite-Based Compiler Middle-End
Our system allows programmers to specify custom rewrite rules in Julia syntax. These rules are then applied on an e-graph representing the code of a method. Once the rules are applied and the e-graph is saturated or a timeout has been reached, the compiler will try to extract the optimal code embedded within that e-graph.
@@
Rules have a left-hand side and a right-hand side. Each side consists of a regular Julia expression, i.e., of a tree of function calls, plus variables and/or literals. Two example rules are and . In these rules sin, cos, ^, +, and translate are the names of the called functions. The ~ prefix identifies pattern variables in a rule. These can match any input values of the expression’s function calls, be it the values of other expressions, variables, or literals. The :: notation is used to specify the types of values for which the rule can be applied. If no type is specified, values of any type can be matched. A rewrite rule can be applied whenever the pattern of function calls, literals, and variables on its left-hand side matches an IR code fragment.
@@
As expressions in Julia source code and the SSA IR produced for them contain the same function calls, it is straightforward to search for such matches in a method’s SSA IR code. In other words, there is no need to have rule authors express equivalences in terms of IR constructs. They can instead express them in regular source code expressions similarly to how those would occur in source code to be optimized. This syntax using regular Julia code expressions, with the addition of unbound, optionally typed variables, covers all rewrites where a single expression tree is replaced with a new expression tree. This syntax does not allow rules matching multi-patterns (Jia et al., 2019; Yang et al., 2021). Such patterns can match multiple expression trees and rewrite them each, potentially reusing variables in multiple output expressions.
@@
As in many existing compiler middle-ends, the optimization based on these rewrite rules/patterns will operate on an intermediate representation (IR) that supports optimizations while still preserving high-level language-specific information (Lattner et al., 2021; Li et al., 2021; swiftlang, 2012). In our case, this is the Julia compiler’s SSA-based IR (Bezanson et al., 2012).
A consideration to be made for pattern rewriting is at what point in the compilation to apply patterns. Concretely, patterns can be matched on IR before or after function calls have been inlined. Before inlining, the current method’s IR still contains function calls as they were written by the programmer. At that stage in the compilation, high-level rewrite rules can typically be matched and applied, i.e., rules involving functions of high-level, abstract, and domain-specific APIs. However, some rewrite opportunities could then remain hidden behind call barriers. This can happen, e.g., when APIs with domain-specific function names wrap more generic computations such as underlying linear algebra computations. If the programmer used the domain-specific wrapper APIs while the rewrite patterns are expressed in terms of the underlying linear algebra operations, those patterns will not be found in the IR code before inlining.
@@
Ideally, when exploring rewrite opportunities, the compiler should hence consider all possible combinations of inlined and not-yet-inlined functions. Support for such an exploration is for future work. Instead, our current prototype implementation performs rewriting optimization only before inlining. We note that many interesting rewrites that are not covered by classical compiler optimization occur before inlining, when high-level API usage has not yet vanished from the code. For this reason also, the example snippets of IR in this paper contain code that has not been inlined yet.
@@ 3.1. From IR to E-Graph and Back Again
Top left: a simple Julia method for function pow that computes x to the power n and that prints a message to illustrate the impact of side effects. Top right: the corresponding Julia IR that is generated if the pow method is generated for two integer arguments. It consists of four basic blocks, one of which includes two phi-functions. Bottom left: the e-graph for the expressions in the method consisting of 11 e-classes, including classes for the phi-functions and for the side effect. Bottom right: the CFG skeleton of the method, which looks like the CFG itself, and in which each node refers to certain e-classes in the e-graph where the control flow depends on expressions modeled by those e-classes.
To apply rewrites on the Julia IR of a method, we first convert it to an e-graph and a CFG skeleton. For this, we leverage methods first introduced in the Cranelift compiler (Fallin, 2023). Fig. 3 shows an example Julia function pow together with its SSA IR, e-graph, and CFG skeleton. The structured control flow present in the original source code has been lowered into four basic blocks in the Julia IR, each terminated by an instruction that jumps to another block or returns.
Each e-node in the e-graph corresponds to an SSA value in the IR. For example, the value %7 in the IR corresponds to the e-node labeled -. The dotted boxes around the e-nodes represent the different e-classes. Since no equalities have been registered yet in this e-graph, each e-class consists of exactly one e-node. In this representation, the phi-nodes are leaf nodes, i.e., their dependencies are not included in the e-graph. This is because an SSA phi function introduces a branch in an expression that cannot be represented in an e-graph directly. The lack of phi-node information implies (i) that rewrites including phi-node dependencies are not supported, and (ii) that the e-graph by itself does not contain sufficient information to reconstruct all control flow of the original code. For (i), an example of a useful rewrite that is not supported is to reduce a phi-node to a value if all possible values in the phi-node are equivalent.
For (ii), it is the CFG skeleton that provides the extra information needed to reconstruct valid IR. To that extent, the CFG skeleton contains all operations responsible for control flow and all operations with side effects, including function calls that have side effects. All operations in the skeleton reference the e-classes they depend on. To visualize these references, we use the roman numerals in Fig. 3. Each roman numeral labels an e-class of which the corresponding SSA value is used in an operation with side effects, or in a control flow statement. For example, since a call to println has side effects, it is stored in the third block of the CFG skeleton, referencing e-class IV.
Cranelift’s SSA IR format, like MLIR (MLI, 2019) and Swift (swiftlang, 2012), uses basic block arguments. These offer an alternative to phi-nodes for representing control-dependent data flow in SSA IRs. Whereas phi-nodes represent such data flow within a basic block, basic block arguments represent such data flow at the basic block level. Even though Julia IR contains phi-nodes, we opt to use block arguments in the CFG skeleton because all information regarding out-of-block uses of expressions that need to be materialized within a block is then explicitly represented within that block itself.
To illustrate, the references to e-classes I, II, V, and VI in the two goto #2 statements in the skeleton are basic block arguments. They are passed as arguments to control flow statements and can be references in the destination block, similarly to how functions have arguments. The relation between basic block arguments and phi-nodes in the original Julia IR is simple. For example, the two phi-nodes in block #2 in the Julia IR correspond to the two basic block arguments passed in the CFG skeleton by blocks #1 and #3. Note that while the e-graph in Fig. 3 contains nodes labeled phi, these are purely labels and do not carry the information contained in SSA phi-functions. The phi-nodes in the e-graph can be interpreted as being the corresponding basic block arguments.
It is the CFG skeleton that allows the conversion back to valid Julia IR. As long as each statement in the CFG skeleton is elaborated, meaning that that statement and all of its dependencies are materialized in the generated IR, the resulting program has the same control flow and side effects as the original program.
@@
It is quite common that part of a method depends on a kwarg, e.g., to perform extra checks or to modify the behavior. See, e.g., the use of the flipkernel kwarg in NNLib.jl (NNl, 2025).
@@
BLAS/LAPACK-like libraries often use characters to indicate the format of an array (transposed or not, upper/lower triangular, etc.). Such checks are then independent of the values calculated beforehand. For example, in the LinearAlgebra.jl standard library (Lin, 2025), the costly operation blks = findall(...) is first computed, after which there is a branch based on uplo, followed by a reuse of blks. Another example is CUDA.jl’s CUSPARSE wrappers: Plenty of API calls (which are a prime use case for e-graph-based rewriting) occur outside of fmt checks, others occur inside such checks (CUD, 2025a).
@@
Continuing the subject of API calls, it is common that certain API uses are conditional on the version of the library. For example, CUDA.jl includes checks such as CUSPARSE.version() v"11.7.2" (CUD, 2025b).
@@
In scalar ‘mathy’ code, this pattern is also common. The absolute cost of operations outside the if is then often not that great, but when it comes to optimizing very high-performance code, it can make a difference. Some examples can be found in the standard float.jl (flo, 2025) and math.jl (mat, 2025) libraries.
@@
Our implementation uses Metatheory.jl (Cheli, 2021b), a Julia EqSat package based on egg (Willsey et al., 2021). At its core, it provides an efficient implementation of the e-graph datastructure and EqSat algorithms, and a system that allows to register new types of e-nodes. In our work, we use this system to implement e-nodes that represent Julia IR statements. Metatheory.jl also offers macros to easily specify rewrite rules with a high-level, declarative syntax. We extended this rule syntax to support type annotations, as will be discussed in Section 3.2. Lastly, Metatheory.jl contains a generic extraction algorithm. However, this algorithm is unsuited for our problem, as extraction needs to take into account the information from the CFG skeleton as well. In Section 3.5 we will discuss our extraction approach.